# Polarized multiwavelength emission from pulsar wind – accretion disk interaction in a transitional millisecond pulsar

### v1

@@ Acknowledgments
Appendix A Observations & data processing
The multiband data presented in this work were acquired by different space-borne and ground-based telescopes (Table 2). In the following, we describe the observations and the procedures adopted to process these data. In the following analyses, all timestamps are referenced to the barycenter of the Solar System using the Jet Propulsion Laboratory planetary development ephemeris DE440 [100] and the position derived from observations with the Very Long Baseline Array [36]: RA = 10h23m47s69 and decl. = +00∘38′40.′′8 (Julian epoch date J2000.0; the uncertainties in both coordinates are 2 as).
A.1 IXPE
The Imaging X-ray Polarimetry Explorer (IXPE), a joint mission between the National Aeronautics and Space Administration (NASA) and the Italian Space Agency (ASI), was launched on December 9, 2021. The observatory consists of three identical telescopes, each equipped with an X-ray mirror module assembly that uses grazing-incidence optics to focus X-rays onto a detector unit (DU). Each DU contains a polarization-sensitive gas-pixel detector (GPD) that measures linear X-ray polarization by imaging the tracks of photoelectrons produced by incoming X-rays absorbed in the special fill gas [12, 39]. The GPDs are capable of recording the arrival time of X-ray photons with a time resolution of 100 s (for information on IXPE, see [111, 127]).
IXPE observed J1023 during two sessions: the first from 2024 May 29 at 12:32 UTC to 2024 June 2 at 18:12 UTC (MJD 60459.52 – 60463.76), and the second from 2024 June 3 at 21:15 UTC to 2024 June 14 at 14:58 UTC (MJD 60464.88 – 60475.62). The total exposure time per DU was 675 ks. Data processing was performed using the ixpeobssim package [11]. Both spectral and spectro-polarimetric analyses were conducted with HEASoft and the standard FTOOLS, utilizing the latest files in the IXPE Calibration Database (XRT version released on 2024 August 20, GPD version released on 2024 February 28).
Instrumental background in the photon event lists was reduced by applying the rejection algorithm by [40]777https://github.com/aledimarco/IXPE-background. Given the source faintness, background subtraction was also performed. Source and background photons were selected from the cleaned level-2 event data. Source photons were selected from a circular region with a 60′′ radius centered on the source, corresponding to roughly the 95% encircled energy fraction for the point spread function of an on-axis source. The background was selected from an annular region with inner and outer radii of 150′′ and 300′′, respectively (Fig. 7). Even after background rejection, the background signal dominated above 6 keV. Consequently, the polarimetric analysis was restricted to the 2–6 keV energy band.
A.2 NICER
The Neutron Star Interior Composition Explorer Mission (NICER; [52]) is a non-imaging X-ray observatory on board the International Space Station, with a field of view of 30 arcmin. The X-ray timing instrument of NICER consists of 56 co-aligned X-ray concentrator optics paired with silicon-drift detectors, 52 of which are currently operational. NICER is sensitive to photon energies in the 0.2–12 keV range, provides an effective collecting area of 1900 cm2 at 1.5 keV, and offers a time resolution below 100 ns. During the IXPE campaign, we conducted a couple of NICER observations of J1023 as part of the director’s discretionary time (see Table 2 for a journal of the observations). These observations were aimed at guiding the selection of high and low-mode episodes in the IXPE dataset and were performed during orbit night-time, ensuring high-quality data unaffected by the optical light leak caused by a damage of the thermal shield on one detector sustained in May 2023888See https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/light-leak-overview/.. We reprocessed the data using the NICER Data Analysis Software (NICERDAS) version 13 distributed along with the HEASoft software package and the calibration products released on 2024 February 6. We calibrated and screened the raw unfiltered data and created level-2 cleaned event files using the nicerl2 pipeline with default settings. We extracted time series over the 0.5–10 keV energy range, binned at 10 s, using the nicerl3-lc pipeline. To estimate the background level, we applied the SCORPEON model999https://heasarc.gsfc.nasa.gov/docs/nicer/analysis_threads/scorpeon-overview/ along with the latest geomagnetic data available at the time of analysis (see the yellow dashed lines in Fig. 9, bottom panel).
A.3 Swift
We performed a couple of observations with the Neil Gehrels Swift Observatory (Swift; [51]) simultaneously with the VLA and VLT observations on 2024 June 4–5 (Table 2). During these observations, Swift utilized the X-ray telescope (XRT; [21]) in photon counting mode (readout time of 2.51 s) and the Ultra-Violet/Optical Telescope (UVOT; [106]) in event mode (readout time of 11 ms) with the UVM2 filter in place (central wavelength of 226 nm; full width at half maximum of 52.7 nm). Data were processed using the calibration files released on 2024 May 22 for XRT and on 2024 February 1 for UVOT.
After performing a standard screening of the XRT data, we collected source photons using a circular aperture with a radius of 47.2′′ centered on J1023 and estimated the background level using an annular region with inner and outer radii of 94.4′′ and 188.8′′, respectively. We then extracted background-subtracted time series in the 0.5–10 keV range, binned at 50 s (Fig. 9).
For the UVOT data, we used the coordinator and uvotscreen tools to convert raw coordinates to detector and sky coordinates, filter out hot pixels on the charge-coupled devices, and obtain cleaned event lists. We adopted a circular extraction region of radius 5′′ centered on the position of J1023 and a circular extraction region of radius 10′′ away from the source to collect source and background photons, respectively. We ensured that the extraction region was always centered on the source position in each observation snapshot, and extracted background-subtracted time series binned at 30 s using uvotevtlc (Fig. 9).
A.4 VLT
We observed J1023 with the FOcal Reducer/low dispersion Spectrograph 2 (FORS2; [2]) mounted on the Unit Telescope 1 (Antu) of the 8.2-m Very Large Telescope (VLT) at Paranal Observatory (Chile) in polarimetric mode (Table 2). The observations began on 2024 June 4 at 23:34:37 UTC (MJD 60465.982) and ended on 2024 June 5 at 00:39:23 UTC (MJD 60466.027), and were performed under photometric conditions (seeing 0.3”). A total of 56 images were acquired using the optical filter (; central wavelength of 655 nm; full width at half-maximum of 165 nm), each with an exposure of 20 s, resulting in a total observing time of 3946 s, overheads included. A Wollaston prism was placed in the light path of the instrument, splitting the incident radiation into two orthogonally polarized beams (ordinary and extraordinary). A Wollaston mask was used to prevent these beams from overlapping on the CCD. Moreover, a rotating half-wave plate (HWP) was installed, allowing images to be captured at four distinct angles () relative to the telescope axis: . This step is critical to obtain polarization measurements because combining images taken at multiple (more than 2) angles is necessary to determine the level of linear polarization. Four sets of images were acquired with this setup, one for each angle of the HWP. This process was repeated 14 times to obtain the entire dataset.
Images were processed by subtracting an average bias frame and dividing by a normalized flat field. Aperture photometry was performed using the daophot tool [114], using a 6-pixel aperture. Using Equations 1–2 of [9], we calculated the normalized Stokes parameters and for the linear polarization. These parameters have not been corrected for instrumental contributions to the linear polarization. However, FORS2 routinely observes unpolarized standard stars to monitor instrumental polarization, and this has consistently remained very low (below 0.3%) across all bands over the past decade.
To estimate the degree and angle of linear polarization, we evaluated the parameter for each of the HWP angles, following [41] and references therein (see also [34, 9, 10]). describes the linear polarization along the direction (e.g., ), and it is defined as:
Here and are the extraordinary and ordinary fluxes at a given angle , respectively, and and are the corresponding fluxes for a non-polarized standard star in the field. The relationship between , the polarization degree and the angle is given by:
The parameters and were determined through a two-step process: 1) maximizing the Gaussian likelihood function to provide an initial estimate for and ; 2) refining the estimates from 1) using a Markov Chain Monte Carlo (MCMC) procedure [67] (for further details on the algorithm, see [9, 10]). The best-fit values were obtained as the median of the marginal posterior distributions from the MCMC procedure. The uncertainties were estimated using the 16th to 84th percentiles of these distributions.
@@ A.4 VLT
Since is calculated by normalizing with a non-polarized star in the field, the resulting value for the polarization degree obtained after fitting the curve is already corrected for instrumental effects. However, the polarization angle still requires correction using the Stokes parameters of a polarized standard star. To achieve this correction, we observed the star Vela 1-95 and measured its Stokes parameters to accurately determine its polarization angle. This measured angle was then compared to the tabulated value, resulting in a correction of applied to the polarization angle of J1023. Finally, to correct for the interstellar contribution, instead of using only one non-polarized star, we used a group of reference stars in the same field as the target in Equation 2, assuming these stars are non-polarized. This method allows the polarization degree of the target to be corrected for both instrumental and interstellar contributions. It is crucial to ensure that the chosen reference stars all exhibit the same polarization degree, which would be a good indicator of their intrinsic non-polarized nature (see e.g., [9]). Since the field of J1023 is relatively empty of optical sources, we could select only four comparison stars. However, these stars clustered very well together in the plane, allowing us to determine the instrumental and interstellar polarization quite precisely in each image. We note that, as reported by [10], the maximum interstellar contribution to the linear polarization degree of J1023 is quite low (), and comparable to the polarization measured for the instrument configuration.
A.5 VLA
@@
We acquired radio observations of J1023 – simultaneous with the IXPE and VLT monitoring – using the Karl G. Jansky Very Large Array (VLA; Project ID 24A-476) radio interferometer. Our full observations, including both target and calibrator scans, began on 2024 June 4 at 23:00:00 UTC (MJD 60465.958) and ended on 2024 June 5 at 02:59:20 UTC (MJD 60466.125). The source underwent a bright multi-wavelength flare starting at 00:36:00 UTC on 2024 June 5 (MJD 60466.025); in this work, we only include the analysis of the pre-flare dataset (i.e., 2024 June 4 at 23:00:00 to 2024 June 5 at 00:36:00, with 1.2 hr on source).
We observed in C-band (central frequency of 6 GHz) using the 3-bit sampler, which breaks the total un-flagged bandwidth of 4096 MHz into 32 spectral windows, each comprised of 64 2 MHz frequency channels. We used 3C286 as the flux and bandpass calibrator, J10240052 as the nearby complex gain calibrator (i.e., the secondary), and the unpolarized source, J14072827, as the leakage calibrator. The target monitoring of J1023 was broken into 9-min scans, with two 1-min secondary scans bookended each target scan.
We flagged and calibrated the parallel-hand visibilities with the Common Astronomy Software Applications package (CASA) VLA pipeline [26]. After the pipeline processing was completed, we inspected the visibilities, removing residual radio frequency interference (RFI) through a combination of manual editing and CASA auto-flagging algorithms (i.e., rflag and tfcrop). After the parallel-hand calibration (which is only sufficient for Stokes ) we calibrated the cross-hand visibilities following the procedure outlined in the VLA tutorials101010The tutorial can be found here: https://casaguides.nrao.edu/index.php/CASA_Guides:Polarization_Calibration_based_on_CASA_pipeline_standard_reduction:_The_radio_galaxy_3C75-CASA6.5.4.

### v2

@@ 2 The multi-wavelength dataset
2.1 IXPE Observations
IXPE observed J1023 during two sessions: May 29–June 2 and June 3–June 14, 2024 (MJD 60459.52–60463.76; MJD 60464.88–60475.62). The total exposure time per detector unit (DU) was 675 ks. To reduce instrumental background, we applied the rejection algorithm by A. Di Marco et al. (2023)111https://github.com/aledimarco/IXPE-background. Source photons were extracted from a 60′′-radius circular region centered on the source, and background photons were taken from an annular region with inner and outer radii of 150′′ and 300′′, respectively. We classified the emission modes based on count-rate thresholds from previous studies (e.g., S. Bogdanov et al. 2015; F. Coti Zelati et al. 2018) and refined them with a statistical analysis of the IXPE data. A detailed description of the selection process is provided in Appendix A. Due to background dominance above 6 keV, we restricted the polarimetric analysis to the 2–6 keV band. Data processing was performed using the ixpeobssim package (L. Baldini et al., 2022). Spectral and spectro-polarimetric analyses were conducted with HEASoft and FTOOLS using the latest calibration files.
2.2 NICER Observations
2.3 Swift Observations
2.4 VLT Observations
We observed J1023 with the FOcal Reducer/low dispersion Spectrograph 2 (FORS2) on the VLT in polarimetric mode (Table 1). The observations were performed from 2024 June 4 at 23:34:37 UTC to 2024 June 5 at 00:39:23 UTC, under photometric conditions (seeing 0.3”). A total of 56 images were acquired using the filter ( band; central wavelength 655 nm; FWHM 165 nm), each with a 20 s exposure, totaling 3946 s. These observations are part of a longer dataset, which includes an additional 3796 s of observations with the same setup, performed after the conclusion of the observations presented here (Baglio, M. C. et al., 2025). A Wollaston prism and rotating half-wave plate (HWP) allowed images at four angles , . Four sets of images were acquired for each angle, repeated 14 times.
The images were bias-subtracted and flat-fielded. Aperture photometry was performed using the daophot tool (P. B. Stetson, 1987) with a 6-pixel aperture. Using Equations 1–2 of M. C. Baglio et al. (2020), we calculated the normalized Stokes parameters and . To estimate the degree and angle of linear polarization, we evaluated the parameter for each of the HWP angles, following S. di Serego Alighieri (1998) and references therein (see also S. Covino et al. 1999; M. C. Baglio et al. 2020, 2023). This is linked to the the polarization degree and the angle by the formula:
We initially estimated the polarization degree and angle by maximizing a Gaussian likelihood function and then refined these estimates using a Markov Chain Monte Carlo (MCMC) procedure (D. W. Hogg & D. Foreman-Mackey, 2018) (for further details on the algorithm, see M. C. Baglio et al. 2020, 2023). The polarization angle was calibrated using observations of the polarized standard star Vela 1-95, resulting in a correction of .
@@ 2.4 VLT Observations
Instead of relying on a single unpolarized star, we used several field stars (assumed to be intrinsically unpolarized) to correct for both instrumental and interstellar polarization. It is crucial to ensure that the chosen reference stars exhibit similar polarization degrees, which would be a good indicator of their intrinsic non-polarized nature (see e.g., M. C. Baglio et al. 2020). Since the field of J1023 is relatively empty of optical sources, we could select only four comparison stars. However, these stars clustered very well together in the plane, allowing us to determine the instrumental and interstellar polarization quite precisely in each image. As reported by M. C. Baglio et al. (2023), the maximum interstellar contribution to the linear polarization degree of J1023 is quite low (), and comparable to the polarization measured for the instrument configuration.
2.5 VLA Observations
@@
We obtained simultaneous VLA radio observations of J1023 with IXPE and VLT from 2024 June 4 at 23:00:00 UTC to June 5 at 02:59:20 UTC (Project ID 24A-476). The source underwent a bright multi-wavelength flare starting at 00:36:00 UTC on June 5. We include here only the pre-flare data, yielding about 1.2 hr on-source. The full dataset is presented by Baglio, M. C. et al. (2025).